Concord Company makes radios that sell for $70 each. For the coming year, management expects fixed costs to total $210,000 and variable costs to be $35 per unit.

Compute the break-even point in dollars using the contribution margin (CM) ratio. (Round answer to 0 decimal places, e.g. 1,225.)

Break-even point

=

Compute the margin of safety ratio assuming actual sales are $800,000. (Round margin of safety ratio to 2 decimal places, e.g. 10.51.)

Margin of safety

Compute the sales dollars required to earn net income of $875,000.

Required sales

$
Answer #1
Unit Contribution margin 35 =70-35
Contribution margin (CM) ratio 50% =35/70
1
Fixed costs 210000
Divide by Contribution margin (CM) ratio 50%
Break-even point 420000
2
Margin of safety sales 380000 =800000-420000
Margin of safety ratio 47.50% =380000/800000
3
Fixed costs 210000
Add: Net income 875000
Required Contribution margin 1085000
Divide by Contribution margin (CM) ratio 50%
Required sales 2170000